OpenStudy (anonymous):

A square has an area of 49^2+84x+36 find the length of a side

OpenStudy (anonymous):

49x^2+84x+36, right?

OpenStudy (anonymous):

in that case, (7x+8)

OpenStudy (anonymous):

no, the correct answer turned out to be 7x+6

OpenStudy (anonymous):

thank you though

OpenStudy (anonymous):

7x+6, right. Sorry. 49=7^2, 36=6^2 Typo. Forgive me!

OpenStudy (anonymous):

and of course 2*7*6=84